# Physical partisan proximity outweighs online ties in predicting US voting outcomes

### v1

@@ Methods
We define partisan exposure of a county $i$ as the relative exposure to people with a certain voting behavior (i.e., Republican or Democratic voters), weighted by either the co-location probability between counties $i$ and $j$ (i.e., Colocation Maps) or the relative probability of friendship on Facebook between $i$ and $j$ (i.e., the Social Connectedness Index). Formally, we define the partisan exposure as:
@@
where $PE_{p}(i)$ is the partisan exposure to either Republicans or Democrats for a county $i$, $N$ is the number of US counties, $v_{p}(j)$ is the percentage of Republican or Democratic voters in the county $j$, and $p_{ij}$ is either the co-location probability between counties $i$ and $j$, or the relative probability of friendship on Facebook between $i$ and $j$. Partisan exposure ranges from 0 (low exposure to either Democrats or Republicans) to 1 (high exposure to either Democrats or Republicans). The sum of partisan exposures to Republicans and Democrats is less than $1$ because we account for null and third-party votes.
@@
Residential and experienced partisan segregation are computed from the residential and either online or offline partisan exposures respectively. Specifically, partisan segregation is computed as the difference between exposure to Republicans and to Democrats. The index is scaled from 0 (indicating exclusive exposure to Democrats) to 1 (indicating exclusive exposure to Republicans). Intermediate values indicate varying levels of partisan mixing. The index is computed as follows:
@@
Swing counties
We define swing counties as counties that experienced different majority outcomes in consecutive presidential elections between 2012 and 2020 (for example, a majority of Republican votes in 2012 followed by a majority of votes for the Democratic candidate in 2016). The number of swing counties according to our definition is $285$. Hence, swing counties can determine political elections with their results. Figs. 3e and 3f show their geographic locations and classification according to the RUCC. Swing counties, as defined in our study, are located in both metropolitan and non-metropolitan areas, mainly concentrated in the Midwest and Northeast regions. We also perform the same analyses defining swing counties as counties that experienced different outcomes between 2008 and 2020, for a total of $455$ counties (see SI Appendix, Fig. S2).
@@
To explore which dimension of partisan exposure better predicts voting patterns in US counties, we separately model physical, online, and residential exposures using spatial regressions, both for Republicans and Democrats. Due to the high correlation between these dimensions, combining them in the same model could impact the interpretation and potentially lead to misleading conclusions. Specifically, we use spatial lag models [40], which are autoregressive models that account for spatial autocorrelation among observations. Specifically, the models are defined as:
@@
where $v$ is the share of either Republican or Democratic votes, $p$ is either Democrats or Republicans, $\rho Wy$ is the autoregressive coefficient, and $PE$ is the partisan exposure to either Democrats or Republicans, $p$. The model is fitted for each dimension $d$, namely offline, online, and residential. To enhance the robustness of the method, we compute k-nearest neighbour spatial weights $W$ using different values of $k$, specifically $5$, $7$, and $10$. This approach is preferred because contiguity-based methods would exclude non-contiguous counties from the analysis. When comparing metropolitan and non-metropolitan areas based on the Rural-Urban Continuum Codes (RUCC), we use ordinary least squares (OLS) regressions. In both types of models, we determine the dimension that better predicts voting patterns based on its ability to explain the variance of the dependent variable (i.e., $R^{2}$).
@@
While we treat the three dimensions separately in the regressions, due to their high collinearity, we use dominance analysis [41] to consider them together and evaluate their relative importance. This method allows us to evaluate the impact of predictors by measuring the contribution of each predictor through designing a series of linear regressions with all possible combinations of predictors to systematically assess and compare each variable’s individual and combined predictive power. Specifically, we perform dominance analysis for both the exposure to Republicans and Democrats, including in the models physical, online, and residential exposure along with the demographic and socioeconomic characteristics previously described. The method provides the relative importance of each predictor as a percentage (Fig. 2). The full model is defined as:
@@
To enhance the robustness and the generalizability of the result related to the relationship between partisan exposure and voting patterns, we model each dimension separately using Elastic Net [47] models with k-fold cross-validation ($k=5$). This allows us to evaluate beta coefficients of the model using a regularization technique.
Second, we predict partisan segregation in metropolitan areas (RUCC 1–3) for each dimension using demographic and socioeconomic characteristics of the US counties. To this aim, we model Gradient Boosting regressions and compute SHAP [48] values to interpret the predictions. Specifically, drawing on game theory, the SHAP framework computes importance scores for each predictor, representing the impact of each variable on the final prediction. Before training the model, we split the dataset into training and test sets using a 70/30 ratio.
Data and code availability
Colocation Maps are available through the Meta Data for Good program by signing a data sharing agreement (https://dataforgood.facebook.com/). The Social Connectedness Index (SCI) is publicly available through the Humanitarian Data Exchange portal (https://data.humdata.org/dataset/social-connectedness-index). Residential partisan exposure is available through the Harvard Dataverse (https://doi.org/10.7910/DVN/A40X5L). The code to reproduce the study’s findings is available at https://github.com/tonmarco.
Acknowledgements
We gratefully acknowledge Michael Bailey for his help with accessing the original version of the Social Connectedness Index. M.To. acknowledges the support of the NRRP MUR program funded by the NextGenerationEU. B.L. acknowledges the support of the PNRR ICSC National Research Centre for High Performance Computing, Big Data and Quantum Computing (CN00000013), under the NRRP MUR program funded by the NextGenerationEU.
Author contribution statement
M.To. collected and analyzed the data, performed the research, and drafted the first version of the manuscript. M.To., B.L., M.Ti. designed the study, interpreted the results, provided critical feedback, helped shape the manuscript, and revised it. All authors approved the final version of the manuscript.

### v2

@@ Methods
We define the partisan exposure of a county $i$ as the relative exposure to people with a certain voting behavior (i.e., Republican or Democratic voters), weighted by either the co-location probability between counties $i$ and $j$ (i.e., Colocation Maps) or the relative probability of friendship on Facebook between $i$ and $j$ (i.e., the Social Connectedness Index). Formally, we define the partisan exposure as:
@@
where $PE_{p}(i)$ is the partisan exposure to either Republicans or Democrats for a county $i$, NN is the number of US counties, $v_{p}(j)$ is the percentage of Republican or Democratic voters in the county $j$, and $p_{ij}$ is either the co-location probability between counties $i$ and $j$, or the relative probability of friendship on Facebook between $i$ and $j$. Partisan exposure ranges from 0 (low exposure to either Democrats or Republicans) to 1 (high exposure to either Democrats or Republicans). The sum of partisan exposures to Republicans and Democrats is less than $1$ because we account for null and third-party votes.
@@
Residential and experienced partisan segregation are computed from the residential and online or offline partisan exposures, respectively. Specifically, partisan segregation is computed as the difference between exposure to Republicans and Democrats. The index ranges from -1 (indicating exclusive exposure to Democrats) to 1 (indicating exclusive exposure to Republicans) and accounts for third-party voters (since the sum of $PE_{rep}(i)$ and $PE_{dem}(i)$ may be less than $1$). Intermediate values indicate varying levels of partisan mixing. The index is computed as follows:
@@
To explore which dimension of partisan exposure better predicts voting patterns in US counties, we separately model physical, online, and residential exposures using spatial regressions, both for Republicans and Democrats. Due to the high correlation between these dimensions (see SI, S2A), combining them in the same model could impact interpretation and potentially lead to misleading conclusions. Specifically, we use spatial lag models [63], which are autoregressive models that account for spatial autocorrelation among observations. Specifically, the models are defined as:
@@
where $v$ is the share of either Republican or Democratic votes, $p$ is either Democrats or Republicans, $\rho Wy$ is the autoregressive coefficient, and $PE$ is the partisan exposure to either Democrats or Republicans, $p$. The model is fitted for each dimension $d$, that is, offline, online and residential. To enhance the robustness of the method, we compute the spatial weights of k-nearest neighbors $W$ using different values of $k$, specifically $5$, 77, and $10$. This approach is preferred because contiguity-based methods would exclude non-contiguous counties from the analysis. When comparing metropolitan and non-metropolitan areas based on the Rural-Urban Continuum Codes (RUCC), we use ordinary least squares (OLS) regressions. In both types of models, we determine the dimension that better predicts voting patterns based on its ability to explain the variance of the dependent variable (i.e., $R^{2}$).
@@
While we treat the three dimensions separately in the regressions, due to their high collinearity, we use the dominance analysis [64] to consider them together and evaluate their relative importance. This method allows us to evaluate the impact of predictors by measuring the contribution of each predictor through designing a series of linear regressions with all possible combinations of predictors to systematically assess and compare the individual and combined predictive power of each variable. Specifically, we perform a dominance analysis for both the exposure to Republicans and Democrats, including physical, online, and residential exposure in the models along with the demographic and socioeconomic characteristics previously described. The method provides the relative importance of each predictor as a percentage (Fig. 3). The full model is defined as:
@@
Logit regressions on survey data
To understand the relationship between partisan exposure and vote choice, we model four logit regressions, differentiating between exposure to Democrats and Republicans for each wave (pre- and post-election). This approach accounts for the complementary nature and negative correlation between questions like "How many of your friends and family are Democrats?" and "How many of your friends and family are Republicans?" The logistic regression models are specified as follows:
$v(i)=\alpha+\beta_{1}PE_{offline}(i)+\beta_{2}PE_{online}(i)+\beta_{n}controls(i)+\epsilon,$ (7)
where $v(i)$ is the respondents’ vote at the 2020 presidential elections, $PE_{offline}$ represents offline partisan exposure based on friends and family voting behavior, and $PE_{online}$ represents online partisan exposure based on Facebook friends’ vote choice. We include controls for respondents’ age, ethnicity, educational attainment, and place of residence (metro or non-metro area), and apply post-stratification weights based on gender, age, education, and ethnicity using the 2020-2024 5-year estimates of the American Community Survey (ACS). Ethnicity and educational attainment are treated as dummy variables, differentiating between the White, Black, and Hispanic categories and between individuals with and without a degree.
@@
To enhance the robustness and generalizability of the result related to the relationship between partisan exposure and voting patterns, we model each dimension using Random Forest and Elastic Net [70] models with k-fold cross-validation ($k=5$). This allows us to evaluate variable importance using tree models and beta coefficients of the model using a regularization technique.
Second, we predict partisan segregation in metropolitan areas (RUCC 1–3) for each dimension using demographic and socioeconomic characteristics of US counties. To this aim, we model Gradient Boosting regressions and compute SHAP [62] values to interpret the predictions. Specifically, drawing on game theory, the SHAP framework computes importance scores for each predictor, representing the impact of each variable on the final prediction. Before training the model, we split the dataset into training and test sets using a 70/30 ratio.
Acknowledgments
We sincerely thank Michael Bailey for his assistance in accessing the original version of the Social Connectedness Index and Ryan Enos for his support with the dataset on residential partisan exposure. We also thank Alberto Acerbi and Mario Quaranta for their comments on an early version of the manuscript.
Funding
M.To. acknowledges the support of the NRRP MUR program funded by the NextGenerationEU. B.L. acknowledges the support of the PNRR ICSC National Research Centre for High Performance Computing, Big Data and Quantum Computing (CN00000013), under the NRRP MUR program funded by the NextGenerationEU. The work of B.L. was partially supported by the following projects: Horizon Europe Programme, grant #101120237-ELIAS and grant #101120763-TANGO. Funded by the European Union. Views and opinions expressed are however those of the author(s) only and do not necessarily reflect those of the European Union or the European Health and Digital Executive Agency (HaDEA). Neither the European Union nor the granting authority can be held responsible for them.
Author contributions statement
M.To. collected and analyzed the data, performed the research, and wrote the first version of the manuscript. M.To., B.L., M.Ti. designed the study, interpreted the results, provided critical feedback, helped shape the manuscript, and revised it. All authors approved the final version of the manuscript.
Data availability
Colocation Maps are available through the Meta Data for Good program by signing a data sharing agreement (https://dataforgood.facebook.com/). The Social Connectedness Index (SCI) is publicly available through the Humanitarian Data Exchange portal (https://data.humdata.org/dataset/social-connectedness-index). Residential partisan exposure is available through Harvard Dataverse (https://doi.org/10.7910/DVN/A40X5L). The ANES 2020-2022 Social Media Survey dataset is available at https://electionstudies.org/data-center/2020-2022-social-media-study/. The code for reproducing the study findings is available at https://github.com/tonmarco.
Competing interests
The authors declare no competing interests.